OpenStudy (anonymous):

Solve using elimination please: -6y = -2x − 10 -3x + 9y = 15

OpenStudy (anonymous):

i think x = 58/105 but im not sure better get it checked

OpenStudy (anonymous):

yeah thats not it haha nice try though :)

OpenStudy (anonymous):

dayum it XD well ur in 12th right??

OpenStudy (anonymous):

a -2, 1 b infinitely no solution c no solution d 4,3

OpenStudy (anonymous):

yes i am

OpenStudy (shrutipande9):

if u have to go by elimination method...what u need to do is...take one equation and make it in the form x=........ and then substitute the x in the next equation

OpenStudy (anonymous):

i honestly have no idea

OpenStudy (javk):

no that's the substitution method @shrutipande9 in the elimination method you either add or subtract the equations to get the answers

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

-6y = -2x − 10 ............(1) -3x + 9y = 15 ..............(2) First, lets put equation (1) into standard form by adding 2x to both sides, giving us: 2x -6y = -10 ...............(3) Do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Good! So our two equations are now: 2x -6y = -10 ...............(3) -3x + 9y = 15 ..............(2) The next step is to make the coefficients of x in both equations equal and opposite. This can be done by multiplying both sides of equation (3) by 3, and by multiplying both sides of equation (2) by 2, giving us: 6x - 18y = -30 ..........(4) -6x +18y = 30 ..........(5) As you can see, equations (4) and (5) are basically the same. Therefore we cannot find a solution by using elimination.

OpenStudy (anonymous):

thank youuuuu

OpenStudy (kropot72):

You're welcome :)

OpenStudy (javk):

theres something wrong with the qyestion...I think both eqns seem to be multiples of the same eqn

OpenStudy (javk):

my bad you already said that :s

OpenStudy (kropot72):

Note that answer choices (a) and (d) are both solutions. There are infinitely many solutions.

OpenStudy (anonymous):

thanks